Mass of NiCl2*6H2O. Please do steps.
Answer #1

Lets see the molecular mass of each

Ni = 58.693 g/mol

Cl = 35.453 g/mol

hence for Cl2 = 2 x 35.453 = 70.906 g/mol

therefore for NiCl2 = 58.693 + 70.906 = 129.599 g/mol

H = 1.0078 g/mol , then for H2 = 2 x 1.0078 = 2.0156

O = 16  g/mol

therefore for H2O = 2.0156 + 16 = 18.0156 g/mol

then for 6H2O = 6 x 18.0156 = 108.0936 g/mol

Now lets add all to get Mass

= 129.599 + 108.0936

= 237.692 g/mol

So our Answer = 237.692 g/mol